A toy car starts at rest on a circular track of radius 2.30 m. If it uniformly increases speed to 1.70 m/s in 3.10 seconds, find the magnitude of the acceleration of the car at the very end of the trip.
Answer #1

v = 1.7 m/s

R = 2.3 m

t = 3.1 sec

Centripetal Acceleration = Tangential Acceleration

a(c) = a(t) = v^2/R = 1.7^2/2.3 = 1.26 m/s^2